OpenStudy (anonymous):

N=0.04-78.84 Number N of state and federal inmates in millions during year x where x > 2002 can be approximate by the following formula. Determine the year in which there were 1.4 million inmates. It was in the yearwhat year

OpenStudy (unklerhaukus):

do you mean N=0.04x-78.84

OpenStudy (anonymous):

the problem is N=0.04-78.84

OpenStudy (unklerhaukus):

there should be an x in your formula

OpenStudy (anonymous):

I am sorry there is a x

OpenStudy (unklerhaukus):

where abouts

OpenStudy (anonymous):

year x

OpenStudy (unklerhaukus):

\[N(x)=?\]

OpenStudy (anonymous):

not sure it says find the year in which there was 1.4 million inmates

OpenStudy (anonymous):

But just help me solve it by N(x)=

OpenStudy (unklerhaukus):

Is one of these you equation? \[N(x)=0.04(x)-78.84\]\[\downarrow\]\[N=0.04x-78.84\]

OpenStudy (anonymous):

it is the second equation

OpenStudy (anonymous):

are ya still there

OpenStudy (anonymous):

N is the symbol used in the question for the number of inmates (in millions). If we want to find out in what year we had 1.4 million inmates all we have to do is substitute N=1.4 into the equation and solve to find x. Can you see what to do from here?

OpenStudy (anonymous):

no i need help

OpenStudy (anonymous):

Ok so can you write out the equation once we substitute N=1.4 in?

OpenStudy (anonymous):

do not know for sure

OpenStudy (anonymous):

If we substitute N=1.4 we will have as our equation: 1.4=0.04x-78.84 Do you understand how and why I did this?

OpenStudy (anonymous):

yes so how do i get the year

OpenStudy (anonymous):

Well x is the number of years after 2002 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So we have to solve the equation: 1.4=0.04x-78.84 to find x. In other words, we have to get x on one side of the equation and everything else on the other side. Do you see what the first step we could take is to do this?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

Well the first thing stopping us from having x on its own on the right hand side of the = is the -78.84. How could we get rid of this?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

Hint: is there something we could add to 0.04x -78.84 to make the right hand side just = 0.04x?

OpenStudy (anonymous):

1.4

OpenStudy (anonymous):

Not quite, how about this: if we add 78.84 to the right hand side then we will have on the right hand side 0.04x +78.84 - 78.84, which will cancel to give 0.04x. Also, what we do to one side we have to do to the other, so on the left hand side we will have 1.4+78.84. Do you understand?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So after we add 78.84 to both sides of the equation what will we have?

OpenStudy (anonymous):

ok so i got 8021 for the left and 78.8 for right

OpenStudy (anonymous):

We are adding 78.84 to both sides of the equation: 1.4=0.04x-78.84 So we will have: 1.4+78.84=0.04x-78.84+78.84 Since -78.84+78.84=0, those terms will cancel on the right and we will be left with: 80.24=0.04x Do you follow my working?

OpenStudy (anonymous):

yes so would it be year of 2004

OpenStudy (anonymous):

How did you get that? I don't think it's the right answer.

OpenStudy (anonymous):

I just throught0.04 meant 2004

OpenStudy (anonymous):

No 0.04 means 4/100. What we want is to get x on its own remember? So if x is being MULTIPLIED by 0.04, what do you think we have to do to get x on its own?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

What is the opposite to multiplying by something?

OpenStudy (anonymous):

dividing

OpenStudy (anonymous):

Exactly, so if our x is multiplied by 0.04, then what should we divide by to get x on its own?

OpenStudy (anonymous):

I do not know

OpenStudy (anonymous):

Let me give a similar example. If 2x=4, then we notice that the x is being multiplied by 2, and so to get x on its own we divide by 2 on both sides to get x=2. Can you apply this method to the question we're doing?

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

I am helping you. We have 80.24=0.04x. x is being multiplied by 0.04, so to get it on its own we need to divide by 0.04 on BOTH SIDES. Doing this we will get: 80.24 divided by 0.04 = x So what do you get for x when you put the left hand side into the calculator?

OpenStudy (anonymous):

2006

OpenStudy (anonymous):

Hooray! Does this answer look sensible, and did you follow the method?

OpenStudy (anonymous):

I am going to go back and look it over but thanks for staying with me and helping me this is my first math class.

OpenStudy (anonymous):

No worries, if you would like my help in any other questions feel free to ask.

OpenStudy (anonymous):

ok thanks